I'm not sure which section this would go under...see, I'm in algebra 3. On my review a question says to write the function
f(x)=x^2 + 18x + 77
in standard form. The answer given is f(x)=(x+9)^2 -4.
I understand the (x+9)^2 part, but for some reason I am stumped about the -4. If someone could please help I'd appreciate it greatly. Thanks so much!

Answer by stanbon(48564)

f(x)=x^2 + 18x + 77
---------
Complete the square on the x-terms:
---
y = x^2 + 18x + (18/2)^2 + 77 - (18/2)^2
y = (x+9)^2 + 77 - 81
y = (x+9)^2 -4
========================
Cheers,
Stan H.